# Code Graph Model (CGM): A Graph-Integrated Large Language Model for Repository-Level Software Engineering Tasks

### v2

@@ 4.2 Training Strategies
In this task, the inputs are subgraphs randomly sampled from large-scale code graphs, with a limited number of nodes. This constraint ensures that the corresponding output code remains below 8,000 tokens, allowing for computational efficiency and manageable context sizes during training. To enhance the meaningfulness of the output code, we employ a hierarchical approach that preserves the inherent dependencies within the code graphs as they are translated into text. Concretely, for higher-level nodes (e.g., REPO and PACKAGE), we position them at the beginning of the output sequence or their respective files to maintain hierarchical consistency. We then utilize the approach from DeepSeek-Coder [14] to perform topological sorting on all file nodes, thereby establishing a structured order for the code content. Lastly, intra-file nodes (e.g., CLASS and FUNCTION) are sorted by line numbers and concatenated within their respective files, culminating in a coherent text sequence that accurately represents the sampled subgraph.
Noisy Fine-tuning: This phase fine-tunes CGM on real-world issue-patch pairs [5], adapting it to practical software debugging and code editing tasks. As displayed in Figure 2(d), the model learns to generate code patches based on two inputs: (i) a subgraph and (ii) a text prompt that indicates the “oracle” files—files that require modification according to the ground-truth patch. The subgraph is constructed by combining the oracle files, their downstream nodes, and one-hop neighbors from the repository-level code graph. To improve model robustness, we intentionally introduce noise into the prompts: 10% include an irrelevant file that doesn’t require modification, while another 10% omit at least one oracle file. This controlled noise exposure helps the model better generalize to real-world scenarios where inputs may be incomplete or contain irrelevant information.
@@ 4.3 The Graph RAG Framework
This section presents our Graph RAG framework, a streamlined extension of CGM designed for automated resolution of real-world repository tasks. The framework consists of four core modules: Rewriter, Retriever, Reranker, and Reader (the proposed CGM). This compact architecture stands in contrast to the SOTA agentless method, which requires ten distinct steps [2].
As illustrated in Figure 2, the framework operates sequentially. First, Rewriter enhances the original issue description to help Retriever identify relevant nodes in the code graph. Retriever then constructs a connected subgraph using both lexical and semantic search techniques. This subgraph serves as input for both Reranker and Reader. Reranker analyzes the subgraph to identify the Top $K$ files likely to be modified. Finally, Reader (CGM) generates the code patch using both the subgraph from Retriever and the selected files from Reranker. Rewriter and Reranker are implemented by prompting the open-source Qwen2.5-72B-instruct [20], while the semantic search in Retriever utilizes the open-source CGE-Large model [60]. In Appendix D, we provide a case study on how CGM solve a specific issue from scratch. Meanwhile, we report the computational costs of our framework, including the cost of code graph construction, in Appendix C.4
@@
Retriever generates a connected subgraph from the code graph for subsequent modules. As shown in Figure 2(b), Extractor nodes (blue nodes) are first identified through string matching with the code elements and keywords extracted earlier. Next, Inferer nodes are located (red nodes) through semantic search, comparing the Inferer’s output with each node’s textual information. These anchor nodes are then expanded to include their one-hop neighbors, capturing local programming dependencies [61]. To ensure connectivity and incorporate upstream information, these expanded nodes are connected to the Root node (REPO in Figure 1). Finally, each FILE node in the subgraph is expanded to include all its internal nodes, aligning with Reranker’s file-level output. The result is a repository-enhanced context subgraph representing the user query, asdenoted by the shaded nodes in Figure 2(b).
Reranker further refines the subgraph generated by Retriever, selecting only the Top $K$ files deemed most likely to be revised. This refinement is necessary because Retriever’s output includes files that may only be referenced and not modified. Reranker operates in two steps: first, it selects $K=10$ files based on the original user query and file names; next, it narrows this selection down to $K=5$ files by individually scoring each one according to how relevant its file skeleton [2] is to the user query. The specific prompt for Reranker can be found in the Appendix F.
@@ 5.1 Repository-Level Issue Fixing
subtableSWE-bench Lite Method LLM Agent % R Rank All DARS Agent [Uncaptioned image]CS-3.5 Yes 47.00 1 6 CGM-SWE-PY [Uncaptioned image]Q2.5-72B No 43.00 2 8 Lingxi NA Yes 42.67 3 10 CodeAct-v2.1 [Uncaptioned image]CS-3.5 Yes 41.67 4 11 PatchKitty-0.9 [Uncaptioned image]CS-3.5 Yes 41.33 5 12 Composio SK [Uncaptioned image]CS-3.5 Yes 41.00 6 14 Agentless-v1.5 [Uncaptioned image]CS-3.5 No 40.67 7 32 Moatless [Uncaptioned image]CS-3.5 Yes 39.00 8 19 Patched.Codes [Uncaptioned image]CS-3.5 Yes 37.00 9 20 CGM-Multi [Uncaptioned image]Q2.5-72B No 36.67 10 23 AppMap [Uncaptioned image]CS-3.5 Yes 36.00 11 24 Agentless Lite [Uncaptioned image]o3-mini No 32.33 13 31 Agentless-v1.5 [Uncaptioned image]GPT-4o No 32.00 14 32 Moatless [Uncaptioned image]DS-V3 Yes 30.67 16 35 SWE-Fixer [Uncaptioned image]Q2.5-72B Yes 24.67 26 51 Lingma SWE-GPT [Uncaptioned image]Q2.5-72B No 22.00 28 57 .
@@
As shown in Table 5.1, our CGM-SWE-PY model achieves a 43% resolution rate on SWE-bench Lite, placing it first among methods utilizing open-source models, second among those that implement open-source methods but use closed-source models, and eighth overall. Notably: (i) When compared to other methods based on open-source models, CGM-SWE-PY outperforms Moatless+DeepSeek-V3 by 12.33% [15], despite DeepSeek-V3’s generally superior performance in various coding benchmarks compared to our LLM decoder Qwen2.5-72B [21]. Furthermore, it exceeds Lingma SWE-GPT by 21%, even though the latter employs carefully curated COT (chain-of-thought) data to boost Qwen2.5-72B’s effectiveness in issue resolution. (ii) In relation to other agentless frameworks, CGM-SWE-PY slightly surpasses Agentless+Claude-3.5-Sonnet by 2.33% and significantly outperforms Agentless+GPT-4o by 11.00%. This achievement is particularly noteworthy given that Agentless leverages a complex ten-step pipeline with more powerful closed-source models, while CGM-SWE-PY operates on a simpler four-step Graph RAG framework. We attribute this success to CGM’s enhanced capacity to interpret both semantic and structural information within repositories. (iii) While the top methods on SWE-bench Lite are entirely closed-source regarding both models and implementations, CGM-SWE-PY’s results are within 10% of these systems. This indicates that CGM-SWE-PY has the potential to compete with leading agent-based methodologies. Compared to other open-sourced model-based methods, CGM significantly narrows the gap between open-source models and closed-source methods in issue-fixing scenarios. (iv) Our multi-task model, CGM-Multi, achieves a resolution rate of 36.67% on SWE-bench Lite, ranking 23rd overall. The relatively lower performance compared to CGM-SWE-PY can be attributed to its broader focus, which encompasses both issue fixing and code completion tasks across Python and Java repositories. (v) We further apply CGM-SWE-PY to a larger Python benchmark—SWE-bench Verified in Table 5.1, where CGM-SWE-PY ranks first again among open weight models, and fifth among methods with open-source system.
@@ 5.2 Repository-Level Code Completion
In this section, we evaluate the CGM’s performance on code completion tasks at the repository level for both Python and Java programming languages. Our evaluation uses two benchmarks: CrossCodeEval and ComplexCodeEval. Concretely, CrossCodeEval focuses on cross-file code completion, while ComplexCodeEval encompasses more intricate tasks, including API recommendations and test case generation. Performance is measured using two metrics: Edit Match (EM) and Edit Similarity (ES), evaluating how similar the generated code is to the ground-truth code. Detailed information regarding the datasets, metrics, and the implementation of baseline models can be found in Appendix C.6. Table 3: Performance comparison on CrossCodeEval and ComplexCodeEval benchmarks. DeepSeek-236B represents DeepSeek-V2.5-236B, Mixtral-123B denotes Mistral-Large-Instruct-2411, and Qwen2.5-72B refers to Qwen2.5-72B-Instruct. Baseline models are evaluated using FIM (Fill-in-Middle) and one-hop expansion. CrossCodeEval ComplexCodeEval Method Java Python Java Python EM ES EM ES EM ES EM ES Mixtral-123B 47.17 82.23 53.92 82.42 37.00 64.81 31.00 62.48 DeepSeek-236B 44.74 83.81 58.54 85.03 36.00 63.08 32.00 60.60 Qwen2.5-72B 37.31 78.78 58.50 81.56 26.00 54.14 28.00 57.12 CGM-Multi-72B 50.21 80.76 61.20 84.30 47.00 78.86 43.00 72.60 Table 3 presents the results for CGM-Multi, which uses Qwen2.5-72B-instruct as its LLM decoder. We compare it with similarly sized large language models, including Mistral-Large-Instruct-123B, DeepSeek-V2.5-236B, and the standalone Qwen2.5-72B-instruct. For all models, context retrieval for code completion is performed by identifying one-hop neighbors of the target file (that requires completion) in the code graph. While CGM-Multi processes the entire subgraph as input, baseline models only receive the textual content from the nodes. Results show that CGM-Multi performs on par with or exceeds other models on CrossCodeEval. More importantly, it greatly outperforms the baseline models on ComplexCodeEval, demonstrating its superior capability in handling complex tasks through comprehensive subgraph analysis. Table 4: Comparison of CGM with RAG variants on CrossCodeEval. Results are reported for Java and Python across multiple base models. Evaluation metrics include EM and ES. CodeLLama-7B DeepSeek-Coder-7B Method Java Python Java Python EM ES EM ES EM ES EM ES NoRAG 20.60 54.50 13.70 44.10 24.20 59.30 19.40 52.50 BM25 23.42 66.13 21.76 69.09 22.49 66.78 23.30 70.84 RepoFuse / / 24.80 71.05 / / 27.92 73.09 RLCoder 26.23 67.61 26.60 72.27 26.09 67.31 30.28 74.42 R2C2 35.60 58.50 23.60 42.90 41.60 64.60 32.70 54.00 CGM-Multi 36.42 75.28 31.03 73.90 41.65 74.76 33.88 71.19 StarCoder-7B Qwen2.5-Coder-7B Method Java Python Java Python EM ES EM ES EM ES EM ES NoRAG 21.60 55.90 17.00 49.50 37.31 78.78 33.63 73.19 BM25 22.16 67.80 22.33 69.60 49.37 82.63 43.15 78.66 RepoFuse / / 24.20 70.82 / / / / RLCoder 24.73 69.08 25.82 72.11 / / / / R2C2 38.10 63.60 30.90 51.90 / / / / CGM-Multi 37.44 73.77 31.00 71.66 51.61 84.62 46.23 82.16 Next, we evaluate CGM against other RAG methods for CrossCodeEval. The comparison includes several established systems: BM25, the default retrieval method in CrossCodeEval [62]; RLcoder [63], which employs reinforcement learning for retrieval optimization; RepoFuse [64], which integrates code graphs during retrieval but converts retrieved code snippets into linear text sequences; and R2C2 [65], which combines retrieved code snippets with Tree-sitter-generated abstract context as the input to the LLM. In our CGM implementation, we still construct input subgraphs by combining target files with their one-hop neighbors. We evaluate these methods using various base models for generation, including CodeLlama-7B, StarCoder-7B, DeepSeek-Coder-7B, and Qwen2.5-Coder-7B. This diverse set of comparison methods enables a comprehensive evaluation of CGM’s effectiveness in long-context retrieval and understanding. As shown in Table 4, CGM typically outperforms other RAG methods, regardless of the base model used, suggesting that graph-based context retrieval is more effective for code completion tasks. Moreover, CGM’s superiority over RepoFuse, which also uses code graphs for retrieval, can be attributed to CGM’s explicit integration of structural information within the subgraph, whereas RepoFuse flattens node context into text sequences, obscuring the explicit dependencies among code entities.
@@ B.2 Handling of Complex Dependiences
During the construction of code graph, we explicitly address both dynamic calls and multiple inheritance in the following way. Dynamic Calls: We employ a conservative resolution approach following the over-approximation principle [66]. When encountering base class method calls (e.g., Base.method()), we include all possible overriding implementations from subclasses in the calls set. This ensures we don’t miss any potential execution paths. Multiple Inheritance: We utilize the Class Hierarchy Analysis (CHA) algorithm [67] to properly handle inheritance relationships, including cases where classes inherit from multiple parent classes.

### v3

@@ 4.2 Training Strategies
In this task, the inputs are subgraphs randomly sampled from large-scale code graphs, with a limited number of nodes. This constraint ensures that the corresponding output code remains below 8,000 tokens, allowing for computational efficiency and manageable context sizes during training. To enhance the meaningfulness of the output code, we employ a hierarchical approach that preserves the inherent dependencies within the code graphs as they are translated into text. Concretely, for higher-level nodes (e.g., REPO and PACKAGE), we position them at the beginning of the output sequence or their respective files to maintain hierarchical consistency. We then utilize the approach from DeepSeek-Coder (Zhu et al., 2024) to perform topological sorting on all file nodes, thereby establishing a structured order for the code content. Lastly, intra-file nodes (e.g., CLASS and FUNCTION) are sorted by line numbers and concatenated within their respective files, culminating in a coherent text sequence that accurately represents the sampled subgraph.
Noisy Fine-tuning: This phase fine-tunes CGM on real-world issue-patch pairs (Jimenez et al., 2024), adapting it to practical software debugging and code editing tasks. As displayed in Figure 3(d), the model learns to generate code patches based on two inputs: (i) a subgraph and (ii) a text prompt that indicates the “oracle” files—files that require modification according to the ground-truth patch. The subgraph is constructed by combining the oracle files, their downstream nodes, and one-hop neighbors from the repository-level code graph. To improve model robustness, we intentionally introduce noise into the prompts: 10% include an irrelevant file that doesn’t require modification, while another 10% omit at least one oracle file. This controlled noise exposure helps the model better generalize to real-world scenarios where inputs may be incomplete or contain irrelevant information.
@@ 4.3 The Graph RAG Framework
This section presents our Graph RAG framework, a streamlined extension of CGM designed for automated resolution of real-world repository tasks. The framework consists of four core modules: Rewriter, Retriever, Reranker, and Reader (the proposed CGM). This compact architecture stands in contrast to the SOTA agentless method, which requires ten distinct steps (Xia et al., 2024).
As illustrated in Figure 3, the framework operates sequentially. First, Rewriter enhances the original issue description to help Retriever identify relevant nodes in the code graph. Retriever then constructs a connected subgraph using both lexical and semantic search techniques. This subgraph serves as input for both Reranker and Reader. Reranker analyzes the subgraph to identify the Top $K$ files likely to be modified. Finally, Reader (CGM) generates the code patch using both the subgraph from Retriever and the selected files from Reranker. Rewriter and Reranker are implemented by prompting the open-source Qwen2.5-72B-instruct (Yang et al., 2024a), while the semantic search in Retriever utilizes the open-source CGE-Large model (CodeFuse, 2024). In Appendix D, we provide a case study on how CGM solve a specific issue from scratch. Meanwhile, we report the computational costs of our framework, including the cost of code graph construction, in Appendix C.4
@@
Retriever generates a connected subgraph from the code graph for subsequent modules. As shown in Figure 3(b), Extractor nodes (blue nodes) are first identified through string matching with the code elements and keywords extracted earlier. Next, Inferer nodes are located (red nodes) through semantic search, comparing the Inferer’s output with each node’s textual information. These anchor nodes are then expanded to include their one-hop neighbors, capturing local programming dependencies (Pan et al., 2024b). To ensure connectivity and incorporate upstream information, these expanded nodes are connected to the Root node (REPO in Figure 2). Finally, each FILE node in the subgraph is expanded to include all its internal nodes, aligning with Reranker’s file-level output. The result is a repository-enhanced context subgraph representing the user query, asdenoted by the shaded nodes in Figure 3(b).
Reranker further refines the subgraph generated by Retriever, selecting only the Top $K$ files deemed most likely to be revised. This refinement is necessary because Retriever’s output includes files that may only be referenced and not modified. Reranker operates in two steps: first, it selects $K=10$ files based on the original user query and file names; next, it narrows this selection down to $K=5$ files by individually scoring each one according to how relevant its file skeleton (Xia et al., 2024) is to the user query. The specific prompt for Reranker can be found in the Appendix F.
@@ 5.1 Repository-Level Issue Fixing
subtableSWE-bench Lite Method LLM Agent % R Rank All DARS Agent [Uncaptioned image]CS-3.5 Yes 47.00 1 6 CGM-SWE-PY [Uncaptioned image]Q2.5-72B No 43.00 2 8 Lingxi [Uncaptioned image]CS-3.5 Yes 42.67 3 10 CodeAct-v2.1 [Uncaptioned image]CS-3.5 Yes 41.67 4 11 PatchKitty-0.9 [Uncaptioned image]CS-3.5 Yes 41.33 5 12 Composio SK [Uncaptioned image]CS-3.5 Yes 41.00 6 14 Agentless-v1.5 [Uncaptioned image]CS-3.5 No 40.67 7 32 Moatless [Uncaptioned image]CS-3.5 Yes 39.00 8 19 Patched.Codes [Uncaptioned image]CS-3.5 Yes 37.00 9 20 CGM-Multi [Uncaptioned image]Q2.5-72B No 36.67 10 23 AppMap [Uncaptioned image]CS-3.5 Yes 36.00 11 24 Agentless Lite [Uncaptioned image]o3-mini No 32.33 13 31 Agentless-v1.5 [Uncaptioned image]GPT-4o No 32.00 14 32 Moatless [Uncaptioned image]DS-V3 Yes 30.67 16 35 SWE-Fixer [Uncaptioned image]Q2.5-72B Yes 24.67 26 51 Lingma SWE-GPT [Uncaptioned image]Q2.5-72B No 22.00 28 57 .
@@
As shown in Table 5.1, our CGM-SWE-PY model achieves a 43% resolution rate on SWE-bench Lite, placing it first among methods utilizing open-source models, second among those that implement open-source methods but use closed-source models, and eighth overall. Notably: (i) When compared to other methods based on open-source models, CGM-SWE-PY outperforms Moatless+DeepSeek-V3 by 12.33% (Aor, 2024), despite DeepSeek-V3’s generally superior performance in various coding benchmarks compared to our LLM decoder Qwen2.5-72B (Liu et al., 2024a). Furthermore, it exceeds Lingma SWE-GPT by 21%, even though the latter employs carefully curated COT (chain-of-thought) data to boost Qwen2.5-72B’s effectiveness in issue resolution. (ii) In relation to other agentless frameworks, CGM-SWE-PY slightly surpasses Agentless+Claude-3.5-Sonnet by 2.33% and significantly outperforms Agentless+GPT-4o by 11.00%. This achievement is particularly noteworthy given that Agentless leverages a complex ten-step pipeline with more powerful closed-source models, while CGM-SWE-PY operates on a simpler four-step Graph RAG framework. We attribute this success to CGM’s enhanced capacity to interpret both semantic and structural information within repositories. (iii) While the top methods on SWE-bench Lite are entirely closed-source regarding both models and implementations, CGM-SWE-PY’s results are within 10% of these systems. This indicates that CGM-SWE-PY has the potential to compete with leading agent-based methodologies. Compared to other open-sourced model-based methods, CGM significantly narrows the gap between open-source models and closed-source methods in issue-fixing scenarios. (iv) Our multi-task model, CGM-Multi, achieves a resolution rate of 36.67% on SWE-bench Lite, ranking 23rd overall. The relatively lower performance compared to CGM-SWE-PY can be attributed to its broader focus, which encompasses both issue fixing and code completion tasks across Python and Java repositories. (v) We further apply CGM-SWE-PY to a larger Python benchmark—SWE-bench Verified in Table 5.1, where CGM-SWE-PY ranks first again among open weight models, and fifth among methods with open-source system.
@@ 5.2 Repository-Level Code Completion
In this section, we evaluate the CGM’s performance on code completion tasks at the repository level for both Python and Java programming languages. Our evaluation uses two benchmarks: CrossCodeEval and ComplexCodeEval. Concretely, CrossCodeEval focuses on cross-file code completion, while ComplexCodeEval encompasses more intricate tasks, including API recommendations and test case generation. Performance is measured using two metrics: Edit Match (EM) and Edit Similarity (ES), evaluating how similar the generated code is to the ground-truth code. Detailed information regarding the datasets, metrics, and the implementation of baseline models can be found in Appendix C.6. Table 3: Performance comparison on CrossCodeEval and ComplexCodeEval benchmarks. DeepSeek-236B represents DeepSeek-V2.5-236B, Mixtral-123B denotes Mistral-Large-Instruct-2411, and Qwen2.5-72B refers to Qwen2.5-72B-Instruct. Baseline models are evaluated using FIM (Fill-in-Middle) and one-hop expansion. CrossCodeEval ComplexCodeEval Method Java Python Java Python EM ES EM ES EM ES EM ES Mixtral-123B 47.17 82.23 53.92 82.42 37.00 64.81 31.00 62.48 DeepSeek-236B 44.74 83.81 58.54 85.03 36.00 63.08 32.00 60.60 Qwen2.5-72B 37.31 78.78 58.50 81.56 26.00 54.14 28.00 57.12 CGM-Multi-72B 50.21 80.76 61.20 84.30 47.00 78.86 43.00 72.60 Table 3 presents the results for CGM-Multi, which uses Qwen2.5-72B-instruct as its LLM decoder. We compare it with similarly sized large language models, including Mistral-Large-Instruct-123B, DeepSeek-V2.5-236B, and the standalone Qwen2.5-72B-instruct. For all models, context retrieval for code completion is performed by identifying one-hop neighbors of the target file (that requires completion) in the code graph. While CGM-Multi processes the entire subgraph as input, baseline models only receive the textual content from the nodes. Results show that CGM-Multi performs on par with or exceeds other models on CrossCodeEval. More importantly, it greatly outperforms the baseline models on ComplexCodeEval, demonstrating its superior capability in handling complex tasks through comprehensive subgraph analysis. Table 4: Comparison of CGM with RAG variants on CrossCodeEval. Results are reported for Java and Python across multiple base models. Evaluation metrics include EM and ES. CodeLLama-7B DeepSeek-Coder-7B Method Java Python Java Python EM ES EM ES EM ES EM ES NoRAG 20.60 54.50 13.70 44.10 24.20 59.30 19.40 52.50 BM25 23.42 66.13 21.76 69.09 22.49 66.78 23.30 70.84 RepoFuse / / 24.80 71.05 / / 27.92 73.09 RLCoder 26.23 67.61 26.60 72.27 26.09 67.31 30.28 74.42 R2C2 35.60 58.50 23.60 42.90 41.60 64.60 32.70 54.00 CGM-Multi 36.42 75.28 31.03 73.90 41.65 74.76 33.88 71.19 StarCoder-7B Qwen2.5-Coder-7B Method Java Python Java Python EM ES EM ES EM ES EM ES NoRAG 21.60 55.90 17.00 49.50 37.31 78.78 33.63 73.19 BM25 22.16 67.80 22.33 69.60 49.37 82.63 43.15 78.66 RepoFuse / / 24.20 70.82 / / / / RLCoder 24.73 69.08 25.82 72.11 / / / / R2C2 38.10 63.60 30.90 51.90 / / / / CGM-Multi 37.44 73.77 31.00 71.66 51.61 84.62 46.23 82.16 Next, we evaluate CGM against other RAG methods for CrossCodeEval. The comparison includes several established systems: BM25, the default retrieval method in CrossCodeEval (Ding et al., 2024); RLcoder (Wang et al., 2024d), which employs reinforcement learning for retrieval optimization; RepoFuse (Liang et al., 2024), which integrates code graphs during retrieval but converts retrieved code snippets into linear text sequences; and R2C2 (Deng et al., 2024), which combines retrieved code snippets with Tree-sitter-generated abstract context as the input to the LLM. In our CGM implementation, we still construct input subgraphs by combining target files with their one-hop neighbors. We evaluate these methods using various base models for generation, including CodeLlama-7B, StarCoder-7B, DeepSeek-Coder-7B, and Qwen2.5-Coder-7B. This diverse set of comparison methods enables a comprehensive evaluation of CGM’s effectiveness in long-context retrieval and understanding. As shown in Table 4, CGM typically outperforms other RAG methods, regardless of the base model used, suggesting that graph-based context retrieval is more effective for code completion tasks. Moreover, CGM’s superiority over RepoFuse, which also uses code graphs for retrieval, can be attributed to CGM’s explicit integration of structural information within the subgraph, whereas RepoFuse flattens node context into text sequences, obscuring the explicit dependencies among code entities.
@@ B.2 Handling of Complex Dependiences
During the construction of code graph, we explicitly address both dynamic calls and multiple inheritance in the following way. Dynamic Calls: We employ a conservative resolution approach following the over-approximation principle (Li et al., 2023b). When encountering base class method calls (e.g., Base.method()), we include all possible overriding implementations from subclasses in the calls set. This ensures we don’t miss any potential execution paths. Multiple Inheritance: We utilize the Class Hierarchy Analysis (CHA) algorithm (Sawin & Rountev, 2011) to properly handle inheritance relationships, including cases where classes inherit from multiple parent classes.